# Continuous-Flow Data-Rate-Aware CNN Inference on FPGA

### v1

@@ I Introduction
Another approach proposed in [37] considers a pixel-after-pixel architecture that is tailored to the data rate of the neural network. In this approach, input data is fetched and then sorted into a buffer structure to allow the calculation of the current output of the layer. However, fetching and sorting the input data results in more overhead and requires more resources.
In this work, we present a new paradigm for the design of CNNs that consists of designing a circuit that processes data in a continuous flow. This means that the data rates and the number of components at the CNN stages are adjusted to the amount of data that is processed. Likewise, idle times are removed by multiplexing parallel streams to obtain a smaller number of streams with valid data during most of the time. This results in a smaller number of resources in the CNN and a higher utilization of them, which is close to 100%. In the context of the state-of-the-art, this approach bridges the gap between stream and unrolled architectures by permitting designs with different degrees of parallelization. The main contributions of this paper are:
@@
  • A thorough analysis of CNN layers from the point of view of achieving continuous flow and reducing idle times, together with the explanation on how to connect multiple layers in a CNN for continuous flow.
  • An analysis of the complexity of the CNN layers depending on the parallelization.
  • The proposed approach has the possibility of designing CNNs with different data rates and has high scalability.
  • An automatic tool to implement continuous-flow CNNs.
  • A comparison to the state-of-the-art showing significant improvements in terms of resources, mainly lookup tables (LUTs).
@@ III-B Continuous Flow with Padding in Convolutional Layers
By contrast, we suggest to use implicit zero padding by setting specific multipliers of the KPU to zero, depending on the column index of the input pixel. Thus, the data order does not change and the input and output are continuous. To illustrate this idea, Fig. 3b shows the data order for a $5\times 5$ feature map with zero-padding $p=1$, where $p$ represents the number of padded zeros at the sides of the feature map. Fig. 3b also highlights the same two sliding windows that were highlighted in the Fig. 3a. However, in Fig. 3b, it can be seen that both sliding windows are now contained inside the feature map, leading to continuous flow at the output.
In general, adding padding to convolutional layers leads to continuous flow at the output as long as $p=\frac{k-1}{2}$. The definition of valid outputs for convolutional layers in (5) is then adapted to
@@ III-D Continuous Flow in Pooling Layers
A typical max-pooling operation is implemented using the pooling processing unit (PPU) for the example of a $2\times 2$ kernel and a stride of $s=2$ is shown in Fig. 5. It uses 2 registers (D), 3 max units that calculate the maximum among their two inputs (MAX), and one line buffer (LD). The circuit works similarly to the KPU, where the line buffer ensures that the correct four pixels are compared to each other accordingly to the sliding window. With the nature of pooling layers to reduce the feature map size, the stride of pooling layers is generally larger than one, which leads to a non-continuous output with a data rate of only $1/s^{2}$. This issue has to be addressed in the CNN implementation with the goal of achieving continuous flow and is discussed later in Section IV-D.
@@ IV-A Running Example
To illustrate the concepts that we present in this section, a running example is used. The parameters of the running example are shown in Table V. It consists of a CNN with five layers, where C1 and C2 are convolutional layers, P1 and P2 are max-pooling layers, and F1 is a fully connected layer.
@@ VI Implementation of Common CNN models
To evaluate our approach, we have implemented
a MobileNetV1 [27] model and analyzed the ResNet18 from the ResNet model family[25]. All models were trained on the ImageNet [42] dataset. The MobileNetV1 architecture heavily relies on the depthwise-separable convolution and consists of one convolutional layer, 13 depthwise-separable convolutional layers, followed by an average pooling layer and a fully connected layer. The number of filters in the convolutional layers is controlled by the factor $\alpha$. The MobileNetV1 model family consists of 4 variations with $\alpha\in\{0.25,0.5,0.75,1.0\}$. The ResNet18 architecture consists of residual layer blocks that are put in sequence, where the suffix 18 in the name stands for the number of layers in the model. In Table VIII, we compare the number of parameters and the resource costs for our proposed architecture (Ours) and for a fully parallel implementation (Ref.). Both implementations always have the same input data rate. The costs for the implementation of the activation functions are not included. Activation functions are stateless, scalar functions that do not have to be adapted to a continuous flow architecture, and thereby are not considered in the comparison. The costs for the layer control circuits, for example, the costs for counters to keep track of the current row and column of the input pixel, are also not included as they only apply on a per-layer basis.
@@ VII Synthesis Results
To further evaluate our approach, Table IX compares our MobileNetV1 to previous implementations on FPGA considering the results provided by Vivado. To implement the CNN, we performed a quantization-aware training to adapt the model to the 8-bit fixed-point format used in the implementation. For the quantization-aware training, we have used the MQUAT framework [^1] MQUAT: www.uni-kassel.de/go/mquat. Using MQUAT, the model has been trained to use 8-bit wide fixed-point formats for weights and activations. Only the output activations of the last layer have been set to 12 bits. The process to implement CNNs has been fully automated by developing a code generator. To implement all arithmetic units, FloPoCo [15] has been used except for the general multipliers, where we used the same concept shown in [37], where a single DSP implements two multiplications. The code generator also provides a bit- and clock-accurate simulator for precise accuracy determination using TensorFlow to allow fast inference under hardware conditions. The code generator performs a worst-case analysis to calculate the width of the adders in the KPUs and FCUs of the design. With the given equations, the code generator can implement layer after layer by calculating the data rate for the next layer based on (8), which then allows the software to generate the right number of FCUs/KPUs needed. Thereby, the code generator derives the number of units, configurations, etc., automatically.
@@
All power results in Table IX are obtained using the Vivado power analysis, with accurate switching activity obtained from a post-synthesis simulation. In general, our approach reaches the highest number of FPS with the lowest latency and the best energy efficiency with just 3.55 mW per inference, whereas the FINN [10] implementation shows a much more efficient use of DSP resources relative to the FPS. Compared to [37] on the same FPGA, the number of DSPs and BRAMs used in the proposed CNN is lower by a small margin. The number of LUTs and registers is almost cut in half. Furthermore, our implementation allows a far higher frequency, which reduces the latency by almost half and almost doubles the throughput. Finally, with the quantization-aware training, we reach a better top-1 accuracy for our implementation.
@@
We also implemented and trained a DNN with two dense layers, which have 16 neurons each, with a final third dense layer with 5 neurons. This DNN was then trained with MQUAT on the jet substructure tagging dataset JSC [19], reaching $75.2\%$ top-1 validation accuracy. We then compared different implementations of the DNN with state-of-the-art fully parallel implementations [39, 5, 6, 32, 8, 4, 21].
@@
Most state-of-the-art implementations do not use DSPs in their designs and are, by design, restricted to only use LUT and FF resources. To make our results more comparable to the state-of-the-art, we also created implementations that use LUT resources for the multiplications instead. The LUT multipliers were generated using the IntMultiplier operator from FloPoCo [16], which implements the state-of-the-art approach shown in [11, 12]. For the synthesis results, we used the xcvu9p-flgb2104-2-i FPGA, a Xilinx Virtex Ultrascale+, as it is the same one used in [51], [4], and [39]. Fig. 13 shows a Pareto plot that compares the throughput with LUT utilization for all our implementations compared to the state-of-the-art that reached at least a validation accuracy of $75\%$ and use less than 5k LUTs. We also limit the maximum frequency to the maximum frequency of the FPGA board, which is 800 MHz, as higher frequencies are not possible to reach in practice. The plot shows how flexible our approach is to reach a certain throughput or resource utilization target, and how the Pareto frontier is extended for lower throughput targets.
@@
The plot also shows that fully parallel implementations become more efficient at some point. This is expected, as our approach takes advantage of the reconfiguration of arithmetic units that become less reconfigurable with higher data rates. In Table X, the resources, throughput, validation accuracy, and latency for all state-of-the-art implementations with at least $75.0\%$ validation accuracy are shown together with all of our implementations. It can be seen that without any DSPs and just $1374$ LUTs, a DNN can be implemented that solves the JSC dataset with $75.2\%$ accuracy.

### v2

@@ I Introduction
Another approach proposed in [36] considers a pixel-after-pixel architecture that is tailored to the data rate of the neural network.
In this approach, input data is fetched and then sorted into a buffer structure to allow the calculation of the current output of the layer. However, fetching and sorting the input data results in more overhead and requires more resources.
In this work, we present a new paradigm for the design of CNN accelerators that consists of an architecture that processes data in a continuous flow. A continuous flow architecture is able to process a continuous stream of data at a predefined input data rate while the arithmetic units are matched such that they never have idle times or a lack of data. This means that every arithmetic unit is computing every clock cycle. Internal and output data rates are derived from the CNN model parameters. The architecture allows to process data close to where the data is produced which keeps buffer sizes low. This results in a higher utilization and thereby a smaller number of resources. In the context of the state-of-the-art, this approach bridges the gap between stream and unrolled architectures by permitting designs with different degrees of parallelization. The main contributions of this paper are:
@@
  • A thorough analysis of CNN layers with respect to achieving continuous flow and minimizing idle times, including the explanation for interconnecting multiple layers for continuous flow, as well as an analysis of layer complexity depending on the parallelization.
@@ III-B Continuous Flow with Padding in Convolutional Layers
By contrast, we suggest to use implicit zero padding by setting specific multipliers of the KPU to zero, depending on the column index of the input pixel. Thus, the data order does not change and the input and output are continuous. To illustrate this idea, Fig. 3b shows the data order for a $5\times 5$ feature map with zero-padding $p=1$, where $p$ represents the number of padded zeros at the sides of the feature map. Fig. 3b also highlights the same two sliding windows that were highlighted in the Fig. 3a. However, in Fig. 3b, it can be seen that both sliding windows are now contained inside the feature map, leading to continuous flow at the output. In general, adding padding to convolutional layers leads to continuous flow at the output as long as $p=\frac{k-1}{2}$. The definition of valid outputs for convolutional layers in (5) is then adapted to
@@ III-D Continuous Flow in Pooling Layers
A typical max-pooling operation is implemented using the pooling processing unit (PPU) for the example of a $2\times 2$ kernel and a stride of $s=2$ is shown in Fig. 5. It uses 2 registers (D), 3 max units that calculate the maximum among their two inputs (MAX), and one line buffer (LD). The circuit works similarly to the KPU, where the line buffer ensures that the correct four pixels are compared to each other accordingly to the sliding window.
With the nature of pooling layers to reduce the feature map size, the stride of pooling layers is generally larger than one, which leads to a non-continuous output with a data rate of only $1/s^{2}$. This issue has to be addressed in the CNN implementation with the goal of achieving continuous flow and is discussed later in Section IV-D.
@@ IV-A Running Example
To illustrate the concepts that we present in this section, a running example is used. The parameters of the running example are shown in Table V.
It consists of a CNN with five layers, where C1 and C2 are convolutional layers, P1 and P2 are max-pooling layers, and F1 is a fully connected layer.
@@ VI Implementation of Common CNN models
To evaluate our approach, we have implemented [^1] https://github.com/numericsgithub/CNN-Flow-Paper a MobileNetV1 [27] model and analyzed the ResNet18 from the ResNet model family[25]. All models were trained on the ImageNet [41] dataset. The MobileNetV1 architecture heavily relies on the depthwise-separable convolution and consists of one convolutional layer, 13 depthwise-separable convolutional layers, followed by an average pooling layer and a fully connected layer. The number of filters in the convolutional layers is controlled by the factor $\alpha$. The MobileNetV1 model family consists of 4 variations with $\alpha\in\{0.25,0.5,0.75,1.0\}$. The ResNet18 architecture consists of residual layer blocks that are put in sequence, where the suffix 18 in the name stands for the number of layers in the model. In Table VIII, we compare the number of parameters and the resource costs for our proposed architecture (Ours) and for a fully parallel implementation (Ref.). Both implementations always have the same input data rate. The costs for the implementation of the activation functions are not included. Activation functions are stateless, scalar functions that do not have to be adapted to a continuous flow architecture, and thereby are not considered in the comparison. The costs for the layer control circuits, for example, the costs for counters to keep track of the current row and column of the input pixel, are also not included as they only apply on a per-layer basis.
@@ VII Synthesis Results
The first experiment, we implemented the MobileNetV1, synthesized it using Vivado and compared it to recent state-of-the-art FPGA implementations. Table IX shows the results. To implement the CNN, we performed a quantization-aware training to adapt the model to the 8-bit fixed-point format used in the implementation. For the quantization-aware training, we have used the MQUAT framework [^2] MQUAT: www.uni-kassel.de/go/mquat. Using MQUAT, the model has been trained to use 8-bit wide fixed-point formats for weights and activations. Only the output activations of the last layer have been set to 12 bits. The process to implement CNNs has been fully automated by developing a code generator. To implement all arithmetic units, FloPoCo [15] has been used except for the general multipliers, where we used the same concept shown in [36], where a single DSP implements two multiplications. The code generator also provides a bit- and clock-accurate simulator for precise accuracy determination using TensorFlow to allow fast inference under hardware conditions. The code generator performs a worst-case analysis to calculate the width of the adders in the KPUs and FCUs of the design. With the given equations, the code generator can implement layer after layer by calculating the data rate for the next layer based on (8), which then allows the software to generate the right number of FCUs/KPUs needed. Thereby, the code generator derives the number of units, configurations, etc., automatically.
@@
All power results in Table IX are obtained using the Vivado power analysis, with accurate switching activity obtained from a post-synthesis simulation. In general, our approach reaches the highest number of FPS with the lowest latency and the best energy efficiency with just 3.55 mJ per inference, whereas the FINN [10] implementation shows a much more efficient use of DSP resources relative to the FPS. Compared to [36] on the same FPGA, the number of DSPs and BRAMs used in the proposed CNN is lower by a small margin. The number of LUTs and registers is almost cut in half. Furthermore, our implementation allows a far higher frequency, which reduces the latency by almost half and almost doubles the throughput. Finally, with the quantization-aware training, we reach a better top-1 accuracy for our implementation.
@@
In the second experiment, we performed an analysis for a wide variety of implementations of different data rates on a well known jet substructure tagging dataset JSC [19], which was used in several previous FPGA-based implementations [38, 5, 6, 32, 8, 21]. These implementations were all fully parallel (i.e., one hardware unit per neuron) and all of them except hls4ml are LUT-based, i.e., do no use any DSP (and can not use them). Our DNN consists of two dense layers, which have 16 neurons each, with a final third dense layer with 5 neurons. This DNN was then trained with MQUAT, reaching $75.2\%$ top-1 validation accuracy.
@@
To make our results comparable to the state-of-the-art, we also created implementations that use LUT resources for the multiplications instead. The LUT multipliers were generated using the IntMultiplier operator from FloPoCo [16], which implements the state-of-the-art approach shown in [11, 12]. For the synthesis results, we used the Xilinx Virtex Ultrascale+ FPGA (xcvu9p-flgb2104-2-i), as it is the same one used in [50], [4], and [38].
The detailed results are given in Table X while Fig. 13 shows a Pareto plot that compares the tradeoff between throughput and LUT utilization. The data rate $r_{0}$ of the previous work (top of Table X) is identical to the number of inputs (16) while our results can be tuned to different data rates. The Pareto plot contains all implementations that reached a validation accuracy of at least $75\%$ and use less than 5k LUTs. We also limit the maximum frequency to the maximum frequency of the FPGA clock tree, which is 800 MHz, as higher frequencies are not possible to reach in practice.
@@
The results shows how flexible our approach is to reach a certain throughput or resource utilization target, and how the Pareto frontier is extended for lower throughput targets. The results also show that fully parallel implementations become more efficient at the highest rate $r_{0}=16$. This is expected, as our approach takes advantage of the reconfiguration of arithmetic units that become less reconfigurable with higher data rates.
In detail, one can observe that NeuraLUT-Assemble [6] and TreeLUT [32] achieve lowest resource costs at a higher throughput for $r_{0}=16$. This is due to the architecture design that is specialized for fully parallel implementations and due to the fact that those architectures do not rely on conventual arithmetic but learn LUT interconnections and truth table contents when training. This allows for highly resource efficient implementations but does not allow a sequential implementation. Our approach, in contrast, allows for highly reconfigurable implementations that allows significant lower cost at lower data rates. This flexibility enables designers to trade off resource utilization against throughput according to application requirements. Additionally, as our approach uses conventional arithmetic, it allows utilizing DSP resources too as we can freely decide how to implement multiplications and additions where other approaches that do use LUT-based layers cannot use DSP resources. Using DSPs, we reach a lower LUT resource utilization than NeuralLUT-Assemble at $r_{0}=2$, without DSPs it is achieved at $r_{0}=1/2$. Using DSPs, we reach a lower LUT resource utilization than NeuralLUT-Assemble at $r_{0}=2$, without DSPs it is achieved at $r_{0}=1/2$. Comparing our approach to state-of-the-art conventional arithmetic designs like hls4ml[21], it can be seen that our approach is superior in resources and throughput.